# PACMANN: Point Adaptive Collocation Method for Artificial Neural Networks

### v1

@@ 3 Results
*Adam 0.16% 0.07% 632
Table 3: Overview of the mean and standard deviation of the test error and the mean runtime for each sampling method for the Allen-Cahn equation example.
The mean and standard deviation of the test error and the mean runtime for each sampling method are given in Table 3. Similarly to the previous Burgers’ equation example, the static uniform grid fails to learn the solution and the static Hammersley grid and random resampling offer a significant improvement in accuracy. Out of all sampling methods considered, PACMANN with the Adam optimizer achieves the lowest test error. The next-best method, RAR-D, attains a lower prediction accuracy at the same computational cost. Table 4 demonstrates that Adam results in the lowest mean and standard deviation of the test error for our method in comparison to the other optimization algorithms considered.
Sampling method $L_{2}$ relative error Mean runtime [s] Hyperparameters
@@
*Gradient ascent 0.46% 0.24% 574 $10^{-8}$ 5
*Nonlinear gradient ascent 0.42% 0.24% 602 $10^{-7}$ 5
*RMSprop 0.29% 0.20% 595 $10^{-6}$ 5
*Momentum 0.36% 0.17% 567 $10^{-7}$ 5
*Adam 0.16% 0.07% 632 $10^{-5}$ 5
*Golden section search 0.37% 0.29% 635 $10^{-7}$ 15
Table 4: Overview of the mean and standard deviation of the test error and the mean runtime achieved by PACMANN for each optimization algorithm listed in Section 2, including the optimal values for the stepsize and the number of steps, for the Allen-Cahn equation example.
@@
Figure 6(a) shows the behavior of the different sampling methods when varying the number of collocation points. We find that PACMANN with the Adam optimizer converges the fastest and outperforms the other sampling methods for nearly all numbers of collocation points considered. Most sampling methods stagnate above 2500 collocation points, except for the static uniform grid, which requires up to 10,000 collocation points to compete with the accuracy of the other sampling methods.
@@
Furthermore, we test the behavior of PACMANN with the different optimization algorithms listed in Section 2 by changing the stepsize and the number of steps hyperparameters. Similarly to the previous Burgers’ equation example, the stepsize hyperparameter has a significant influence on the test error, as demonstrated by Figure 7(a). We note the formation of two groups, which occurs due to the random resampling of points that have been moved outside the domain. When using RMSprop and Adam combined with large stepsizes, all 2500 collocation points leave the domain. As a result, the test error converges toward the error of $0.42\%\pm 0.28\%$ found when using random resampling.
@@
Finally, we also vary the number of steps. Figure 7(b) illustrates that the test error of most optimization algorithms is not significantly impacted by changing the number of steps. Figures 7(a) and 7(b) support our earlier finding that the stepsize plays a dominant role in the test error obtained compared to the limited influence of the number of steps.
@@
In the following examples, only the Adam optimizer is evaluated for PACMANN since it consistently outperforms the other optimization algorithms. Furthermore, based on the observations made in the previous Burgers’ equation example and the current Allen-Cahn equation example, we recommend fixing the number of steps to five, which removes a hyperparameter from PACMANN. We recommend five steps as it takes advantage of the increased accuracy achieved by our method with the Adam optimizer at multiple steps (see Figures 3(b) and 7(b)), while keeping the computational cost low (see Figure 4(b)).
3.3 5D Poisson’s equation
To demonstrate the feasibility of using PACMANN for higher dimensional problems, we consider the Poisson’s equation in five dimensions:
[…]
The exact solution is set as
$v\left(\boldsymbol{x}\right)=\prod_{i=1}^{5}\sin(\pi x_{i}),$ (7)
where $x_{i}$ is the i-th component of $\boldsymbol{x}$. For this example, we sample 750 collocation points and 750 points on the boundary condition. Moreover, the network architecture consists of four hidden layers of 64 neurons each.
Sampling method $L_{2}$ relative error Mean runtime [s]
Mean 1 SD
Uniform grid 17.89% 0.94% 742
Hammersley grid 82.08% 3.23% 734
Random resampling 11.03% 0.69% 772
RAR 56.84% 4.46% 753
RAD 10.07% 0.75% 851
RAR-D 88.30% 1.53% 774
*Adam 5.93% 0.46% 778
Table 5: Overview of the mean and standard deviation of the test error and the mean runtime for each sampling method for the Poisson’s equation example.
The mean and standard deviation of the test error and the mean runtime for each of the sampling methods are given in Table 5. In contrast to the previous examples, not only does the static uniform grid fail to capture the solution, but adaptive methods, such as RAR-D, also fail to improve the prediction accuracy. PACMANN in combination with the Adam optimizer and a stepsize of $10^{-2}$ achieves the lowest mean and standard deviation of the test error. Moreover, we point out the ability of this method to efficiently scale to high-dimensional problems. We find that our method with the Adam optimizer is cheaper at a mean runtime of 778 seconds compared to RAD, the next-best adaptive sampling method, at 851 seconds.
Qualitatively, we observe similar behavior of the different sampling methods when varying the number of collocation points and the resampling period for both the Poisson’s equation example and the following Navier-Stokes equation example, as observed in the Burgers’ equation example. This observation also applies to the behavior of PACMANN when changing the stepsize and the number of steps hyperparameters. Therefore, for the sake of the length of this paper, we do not repeat the analysis of the behavior of the sampling methods considered in this study for the Poisson’s equation and Navier-Stokes equation examples.
3.4 2D Navier-Stokes equation
Finally, we consider an inverse 2D Navier-Stokes equation problem of an incompressible flow past a cylinder discussed by Raissi et al. in [5], given by:
[…]
with $(x,y)\in[1,8]\times[-2,2]$ and $t\in[0,7]$. Here, $u$ is the x-component of the velocity field, and $v$ is the y-component. The pressure is denoted by $p$. The scalar parameter $\lambda_{1}$ scales the convective term, and $\lambda_{2}$ represents the dynamic (shear) viscosity. In this example, we are interested in learning the values of the $\lambda_{1}$ and $\lambda_{2}$ based on a data set created by Raissi et al. [5] containing the values of $u$, $v$, and $p$ determined for a large set of points $(t,x,y)$. The true values of $\lambda_{1}$ and $\lambda_{2}$ are 1 and 0.01, respectively. For this inverse problem, we train the PINN on 7000 randomly selected points from this data set. In addition, we sample 700 collocation points, 200 points on the boundary condition and 100 points on the initial condition. The network architecture consists of six hidden layers containing 50 neurons each.
Sampling method $L_{2}$ relative error Mean runtime [s]
@@
Table 6: Overview of the mean and standard deviation of the test error for $\lambda_{1}$ and $\lambda_{2}$ and the mean runtime for each sampling method for the inverse Navier-Stokes equation example.
@@
The mean and standard deviation of the test error for both $\lambda_{1}$ and $\lambda_{2}$ and the mean training time for each of the sampling methods are given in Table 6. In this inverse problem, PACMANN in combination with Adam at a stepsize of $10^{-2}$ achieves the lowest test error for $\lambda_{1}$ and $\lambda_{2}$ at a slightly higher computational cost compared to the following best method, RAR-D. Furthermore, we note that the non-adaptive sampling methods generally outperform the other adaptive methods considered in this study, both in the mean and standard deviation of the test error.
@@ 4 Conclusions
In this work, we presented PACMANN, a novel adaptive collocation point sampling method for physics-informed neural networks. This approach uses the gradient of the squared residual as guiding information to move collocation points toward areas of large residuals, where the size of the step is determined by the stepsize hyperparameter. Our method can be applied in combination with an optimization algorithm of choice, such as gradient ascent or Adam. Points are moved several times while training is halted, given by the ”number of steps” hyperparameter. We explored the sensitivity of our method to its two hyperparameters and we found that the stepsize has a greater influence than the number of steps on the solution accuracy. Conversely, we found that setting the number of steps to 5 achieves the optimal balance between accuracy and efficiency. We then investigated the accuracy and efficiency of PACMANN in combination with various optimization algorithms and concluded that the Adam optimizer performs the best. Furthermore, we compared the performance of PACMANN to existing state-of-the-art adaptive and nonadaptive collocation approaches, including random resampling, RAR and RAD, and demonstrated that our method achieves state-of-the-art performance in terms of accuracy/efficiency tradeoff for lower dimensional benchmarks (1D time-dependent Burgers’ and Allen-Cahn equations), while outperforming the state-of-the-art for the five-dimensional Poisson equation. Finally, we showed the effectiveness of our approach to solve inverse problems in a 2D Navier-Stokes benchmark. The results of the numerical experiments demonstrate that PACMANN achieves high prediction accuracy and efficiently scales to higher dimensions without introducing significant computational overhead.

### v2

@@ 3 Results
PACMANN-Adam 0.16% 0.07% 632
Table 4: Overview of the mean and standard deviation of the test error and the mean runtime for each sampling method for the Allen-Cahn equation example. The best result in each column is marked in boldface.
PACMANN optimizer $L_{2}$ relative error
Mean
runtime [s]
Hyperparameters
@@
Gradient ascent 0.46% 0.24% 574 $10^{-8}$ 5
Nonlinear gradient ascent 0.42% 0.24% 602 $10^{-7}$ 5
RMSprop 0.29% 0.20% 595 $10^{-6}$ 5
Momentum 0.36% 0.17% 567 $10^{-7}$ 5
Adam 0.16% 0.07% 632 $10^{-5}$ 5
Golden section search 0.37% 0.29% 635 $10^{-7}$ 15
Table 5: Overview of the mean and standard deviation of the test error and the mean runtime achieved by PACMANN for each optimization algorithm listed in Section 2. The best result in each column is marked in boldface. We also include the optimal values for the stepsize and the number of steps, for the Allen-Cahn equation example.
@@
In the following example, we consider the one-dimensional Allen-Cahn equation:
[…]
We choose a diffusion coefficient of $d=0.001$. Similar to the Burgers’ equation problem in the previous section, the number of collocation points is set to 2500, the number of boundary points to 80, and the number of initial condition points to 160. The network architecture used for this example consists of four hidden layers of 64 neurons. To compute the $L_{2}$ relative error, we compare the network prediction to a numerical solution generated using a spectral solver and made available via the DeepXDE library [18].
The mean and standard deviation of the test error and the mean runtime for each sampling method are given in Table 4. As for the Burgers’ equation example, the static uniform grid fails to learn the solution satisfactorily, and the static Hammersley grid and random resampling offer a significant improvement in accuracy. Out of all sampling methods considered, PACMANN in combination with the Adam optimizer and a stepsize of $10^{-5}$ achieves the lowest test error. The next-best method, RAR-D, attains a lower prediction accuracy at the same computational cost. Table 5 demonstrates that Adam results in the lowest mean and standard deviation of the test error for our method in comparison to the other optimization algorithms considered.
Figure 7(a) shows the behavior of the different sampling methods when varying the number of collocation points. We find that PACMANN with the Adam optimizer converges the fastest and outperforms the other sampling methods for nearly all numbers of collocation points considered. Most sampling methods stagnate above 2500 collocation points, except for the static uniform grid, which requires up to 10 000 collocation points to compete with the accuracy of the other sampling methods.
@@
Furthermore, we test the behavior of PACMANN with the different optimization algorithms listed in Section 2 by varying the stepsize and the number of steps hyperparameters. Similarly to the results in Section 3.1 for the Burgers’ equation, the stepsize hyperparameter has a significant influence on the test error, as demonstrated by Figure 8(a). We note, again, the formation of two groups, which occurs due to the random resampling of points that have been moved outside the domain. When using RMSprop and Adam combined with large stepsizes, all 2500 collocation points leave the domain. As a result, the test error converges toward the error of $0.42\%\pm 0.28\%$ found when using random resampling; cf. Table 4.
@@
Finally, we also vary the number of optimization steps for resampling. Figure 8(b) illustrates that the test error of most optimization algorithms is not significantly impacted by changing the number of steps. Figures 8(a) and 8(b) support our earlier finding that the stepsize plays a dominant role in the test error obtained compared to the limited influence of the number of steps.
@@
Based on the observations made in the Burgers’ and Allen-Cahn equation examples, we recommend fixing the number of steps to five, which removes a hyperparameter from PACMANN. Furthermore, we recommend five steps as it takes advantage of the increased accuracy achieved by our method with the Adam optimizer at multiple steps (see Figures 3(b) and 8(b)), while keeping the computational cost low; see Figure 4(b).
3.3 2D Navier-Stokes equation (inverse)
Sampling method $L_{2}$ relative error
Mean
runtime [s]
@@
Table 6: Overview of the mean and standard deviation of the test error for $\lambda_{1}$ and $\lambda_{2}$ and the mean runtime for each sampling method for the inverse Navier-Stokes equation example. The best result in each column is marked in boldface.
Next, we consider an inverse problem based on the two-dimensional Navier-Stokes equation describing the flow of an incompressible fluid past a cylinder discussed by Raissi et al. in [6], given by:
$\left\{\begin{aligned} \frac{\partial\boldsymbol{v}}{\partial t}+\lambda_{1}\boldsymbol{v}\cdot\nabla\boldsymbol{v}&=-\nabla p+\lambda_{2}\nabla^{2}\boldsymbol{v},\quad(x,y)\in[1,8]\times[-2,2],\quad t\in[0,7],\\ \nabla\cdot\boldsymbol{v}&=0.\end{aligned}\right$. (13)
Here, $u$ and $v$ are the $x$- and $y$-components of the velocity field, and $p$ denotes the pressure. The scalar parameter $\lambda_{1}$ scales the convective term, and $\lambda_{2}$ represents the dynamic (shear) viscosity. In this example, we are interested in learning the values of $\lambda_{1}$ and $\lambda_{2}$ based on a data set created by Raissi et al. [6] using a spectral solver. The data set contains the values of $u$, $v$, and $p$ determined for a large set of points $(x,y,t)$. The true values of $\lambda_{1}$ and $\lambda_{2}$ are 1 and 0.01, respectively. For this inverse problem, we train the PINN on 7000 randomly selected points from this data set. In addition, we sample 700 collocation points, 200 points on the boundary condition, and 100 points on the initial condition. The network architecture consists of six hidden layers containing 50 neurons each.
@@
The mean and standard deviation of the test error for both $\lambda_{1}$ and $\lambda_{2}$ and the mean runtime for each of the sampling methods are given in Table 6. In this inverse problem, PACMANN in combination with Adam at a stepsize of $10^{-2}$ achieves the lowest test error for $\lambda_{1}$ and $\lambda_{2}$ at a slightly higher computational cost compared to the second best adaptive method, RAR-D. Furthermore, we note that the non-adaptive sampling methods generally outperform the other adaptive methods considered in this study, both in the mean and standard deviation of the test error.
Qualitatively, we observe similar behavior of the different sampling methods when varying the number of collocation points and the resampling period for the 2D Navier-Stokes equation example and the following examples, as observed in the Burgers’ equation example. This observation also applies to the behavior of PACMANN when changing the stepsize and the number of steps hyperparameters, namely, that the number of steps generally has a smaller impact on the prediction accuracy than the stepsize. Therefore, for the sake of conciseness, we do not repeat the analysis of the behavior of the different sampling methods for this two-dimensional Navier-Stokes equation example and the following examples.
3.4 5D Poisson’s equation
Sampling method $L_{2}$ relative error $H^{1}$ semi-norm
Mean
runtime [s]
Mean 1 SD Mean 1 SD
Uniform grid 40.32% 1.18% 62.58% 1.74% 744
Hammersley grid 82.64% 2.95% 90.48% 2.20% 752
Random resampling 11.47% 1.13% 13.36% 1.66% 751
RAR 62.00% 7.13% 69.97% 6.62% 773
RAD 11.69% 1.78% 13.19% 1.31% 841
RAR-D 89.31% 1.98% 93.83% 2.05% 784
PACMANN-Adam 8.35% 0.54% 10.32% 0.43% 786
Table 7: Overview of the mean and standard deviation of the $L_{2}$ relative error and the $H^{1}$ semi-norm, and the mean runtime for each sampling method for the Poisson’s equation example. The best result in each column is marked in boldface.
In the following problem, we apply PACMANN to the Poisson equation in five dimensions:
[…]
We choose the right-hand side function $f$ based on the manufactured solution
$v\left(\boldsymbol{x}\right)=\prod_{i=1}^{5}\sin(\pi x_{i}),$ (15)
where $x_{i}$ is the $i$-th component of $\boldsymbol{x}$. For this example, we sample 750 collocation points and 750 points for the boundary condition. Moreover, the network architecture consists of four hidden layers of 64 neurons each.
The mean and standard deviation of the $L_{2}$ relative error and the $H^{1}$ semi-norm, and the mean runtime for each of the sampling methods are given in Table 7. In contrast to the previous examples, the adaptive methods RAR and RAR-D fail to improve the prediction accuracy. PACMANN in combination with the Adam optimizer and a stepsize of $10^{-2}$ achieves the lowest mean and standard deviation of the $L_{2}$ relative error and the $H^{1}$ semi-norm. Moreover, we point out the ability of this method to efficiently scale to high-dimensional problems. We find that our method with the Adam optimizer is cheaper at a mean runtime of 786 seconds compared to RAD, the next-best adaptive sampling method, at 841 seconds.
3.5 3D Navier-Stokes equation
Sampling method $L_{2}$ relative error
Mean
runtime [s]
$u$ $v$ $w$
Mean 1 SD Mean 1 SD Mean 1 SD
Uniform grid 1.01% 0.16% 1.20% 0.34% 1.27% 0.41% 1506
Hammersley grid 0.67% 0.10% 0.68% 0.09% 0.48% 0.10% 1503
Random resampling 0.50% 0.09% 0.49% 0.09% 0.37% 0.07% 1467
RAR 0.59% 0.13% 0.56% 0.07% 0.42% 0.05% 1462
RAD 0.49% 0.06% 0.51% 0.06% 0.37% 0.03% 1531
RAR-D 0.65% 0.12% 0.68% 0.10% 0.47% 0.06% 1497
PACMANN-Adam 0.46% 0.04% 0.49% 0.09% 0.33% 0.05% 1524
Table 8: Overview of the mean and standard deviation of the $L_{2}$ relative error for $u$, $v$, and $w$ and the mean runtime for each sampling method for the 3D Navier-Stokes equation example. The best result in each column is marked in boldface.
Sampling method $H^{1}$ semi-norm
Mean
runtime [s]
$u$ $v$ $w$
Mean 1 SD Mean 1 SD Mean 1 SD
Uniform grid 1.30% 0.18% 1.40% 0.19% 1.24% 0.26% 1506
Hammersley grid 0.43% 0.05% 0.45% 0.08% 0.29% 0.05% 1503
Random resampling 0.35% 0.04% 0.35% 0.05% 0.25% 0.03% 1467
RAR 0.43% 0.06% 0.42% 0.04% 0.29% 0.03% 1462
RAD 0.36% 0.04% 0.38% 0.06% 0.24% 0.02% 1531
RAR-D 0.43% 0.07% 0.42% 0.05% 0.27% 0.04% 1497
PACMANN-Adam 0.33% 0.03% 0.35% 0.04% 0.22% 0.04% 1524
Table 9: Overview of the mean and standard deviation of the $H^{1}$ semi-norm for $u$, $v$, and $w$ and the mean runtime for each sampling method for the 3D Navier-Stokes equation example. The best result in each column is marked in boldface.
We next apply PACMANN to the incompressible Navier-Stokes equation in a cube:
$\left\{\begin{aligned} \frac{\partial\boldsymbol{v}}{\partial t}+Re\>\boldsymbol{v}\cdot\nabla\boldsymbol{v}+Re\>\nabla p-\nabla^{2}\boldsymbol{v}&=f(\boldsymbol{x}),\quad(x,y,z)\in[-1,1]^{3},\quad t\in[0,1],\\ \nabla\cdot\boldsymbol{v}&=0.\end{aligned}\right$. (16)
Here, we set the Reynolds number to $Re=1$. Furthermore, we choose the right-hand side function $f$ based on the manufactured solution
$u$ $=\cos(x)\sin(y)\sin(z)e^{-t}$ (17)
$v$ $=\sin(x)\cos(y)\sin(z)e^{-t}$
$w$ $=-2\cos(x)\cos(y)\cos(z)e^{-t}.$
We compare PACMANN against the other sampling methods for the Navier-Stokes problem with Dirichlet boundary conditions and initial conditions prescribed by the manufactured solution. For this problem, we sample 300 collocation points, and we employ 75 points for the boundary conditions and 75 points for the initial conditions. The network architecture used for this problem consists of four hidden layers of 64 neurons.
The mean and standard deviation of the $L_{2}$ relative error and the $H^{1}$ semi-norm for $u,v$, and $w$, and the mean runtime for each of the sampling methods are given in Tables 8 and 9. In this problem, PACMANN in combination with Adam at a stepsize of $10^{-4}$ achieves the lowest errors for $u$, $v$, and $w$ while maintaining a computational cost comparable to that of other adaptive sampling methods.
3.6 Re-entrant corner in a disk
Figure 9: Location of the collocation points (a) before training, (b) mid-training at 25 000 iterations, and (c) after training using PACMANN with Adam for the Laplace equation over a re-entrant corner in a disk example. The color indicates the values of the squared residual.
Sampling method $L_{2}$ relative error $H^{1}$ semi-norm
Mean
runtime [s]
Mean 1 SD Mean 1 SD
Uniform grid 6.32% 1.46% 32.36% 6.50% 514
Hammersley grid 0.47% 0.17% 5.90% 0.91% 539
Random resampling 0.49% 0.18% 5.45% 0.82% 536
RAR 0.58% 0.17% 6.12% 1.00% 538
RAD 0.48% 0.25% 4.46% 1.31% 594
RAR-D 0.96% 0.74% 9.39% 6.27% 543
PACMANN-Adam 0.43% 0.17% 5.00% 1.03% 544
Table 10: Overview of the mean and standard deviation of the $L_{2}$ relative error and the $H^{1}$ semi-norm, and the mean runtime for each sampling method for the Laplace equation over a re-entrant corner in a disk example. The best result in each column is marked in boldface.
Finally, we consider a problem involving a domain with a re-entrant corner, where the solution is expected to be less regular. Specifically, we consider the Laplace equation in polar coordinates over a sector:
$\left\{\begin{aligned} u_{rr}+\frac{1}{r}u_{r}+\frac{1}{r^{2}}u_{\theta\theta}&=0,&\quad&r\in[0,1],\quad\theta\in\left[0,\frac{5\pi}{3}\right],\\ u(r,\theta)&=h(r,\theta),&&(r,\theta)\in\partial\boldsymbol{\Omega}.\end{aligned}\right$. (18)
The exact solution is used for the boundary condition $h(r,\theta)$ and is given by
$h(r,\theta)=r^{\frac{3}{5}}\sin\left(\frac{3}{5}\theta\right),$
which has an algebraic singularity at the origin; cf. [36, pp. 110–116]. In this example, we sample 75 collocation points and we set the number of points for the boundary condition to 75. The network architecture used for this problem consists of four hidden layers of 64 neurons.
The mean and standard deviation of the $L_{2}$ relative error and the $H^{1}$ semi-norm, and the mean runtime for each of the sampling methods are given in Table 10. In this example, PACMANN in combination with the Adam optimizer at a stepsize of $10^{-2}$ achieves the lowest $L_{2}$ relative error at a computational cost comparable to that of the adaptive sampling methods RAR and RAR-D, while RAD achieves the lowest $H^{1}$ semi-norm. Furthermore, we observe a collective movement of points toward the origin, where the algebraic singularity is located; see Figure 9.
@@ 4 Conclusions
In this work, we presented the Point Adaptive Collocation Method for Artificial Neural Networks (PACMANN), a novel adaptive collocation point sampling method for physics-informed neural networks. This approach uses the gradient of the physics-informed loss terms, that is, of the squared residual, as guiding information to move collocation points toward areas of large residuals. The problem of moving points is formulated as a maximization problem, which can be approached using an optimization algorithm of choice, such as gradient ascent or Adam. Points are moved several times while training is halted. Our approach can be tuned using three additional hyperparameters, namely the resampling period, the size of the step used to move the collocation points, and the number of times that points are moved while training is halted. While this work demonstrates PACMANN for PINNs, we note that the method can also be applied to other collocation-based approaches.
We studied the sensitivity of our method to these hyperparameters, and we observed that the stepsize has a particularly large impact on the solution accuracy. Conversely, we found that five iteration steps are sufficient to achieve a good balance between accuracy and efficiency. We then investigated the accuracy and efficiency of PACMANN in combination with various optimization algorithms and concluded that the Adam optimizer performs the best.
Furthermore, we compared the performance of PACMANN to existing state-of-the-art adaptive and non-adaptive collocation approaches, including random resampling, RAR, and RAD, and demonstrated that our method achieves state-of-the-art performance in terms of the accuracy/efficiency tradeoff for lower-dimensional benchmarks, while outperforming the state-of-the-art for high-dimensional problems and the case of an irregular domain. In addition, we showed the effectiveness of our approach in solving inverse problems.
In particular, the results of the numerical experiments demonstrate that PACMANN achieves high prediction accuracy and efficiently scales to higher dimensions without introducing significant computational overhead.
Acknowledgments
The authors acknowledge the use of computational resources of the DelftBlue supercomputer, provided by Delft High Performance Computing Centre (https://www.tudelft.nl/dhpc).
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT (OpenAI) in order to improve the readability and language of the manuscript. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.